HOW DO YOU SOLVE THIS?
X-1 - 3X-2 = 7-2X
___ ____ ____
2 4 6

Answer by stanbon(75887)

(X-1)/2 - (3X-2)/4 = (7-2X)/6
Multiply thru by 12 to get:
6(x-1)-3(3x-2) = 2(7-2x)
6x-6-9x+6 = 14-4x
x=14
x=14
Cheers,
Stan H.